alones:

To find the value of y in this proportion you must divide 54 by

alones:

Oh damn..do equations even work?

alones:

\[\frac{ 2 }{ 3 }=\frac{ x }{ 6 }=\frac{ 9 }{ y }\]

celticcat:

by x 2/3 = x/6 so x = 4 so to find y yu divide 54 by 4